OpenStudy (anonymous):

Can someone help me finish this problem I took it as far as I could... If x=9sinθ, use trigonometric substitution to write √(81-x^2) as a trigonometric function of θ, where -pi/2<θ

OpenStudy (anonymous):

So far I have √(81-9sinθ^2) 81√(1-sinθ^2) 1-sin^2θ=1/cscθ ?

OpenStudy (anonymous):

or maybe It is 9√(1-sinθ^2)...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your second post was right

OpenStudy (anonymous):

√(81-81sinθ^2) <--- (9sinθ)^2 = 81sinθ^2 = 9√(1-sinθ^2)

OpenStudy (anonymous):

next you need to figure out which trig substitution to use

OpenStudy (anonymous):

Thank you! would the answer be 9cscθ then?

OpenStudy (anonymous):

\[\sin ^{2}\theta+\cos ^{2}\theta = 1\] is a where you want to start thinking about your substitution

OpenStudy (anonymous):

the answer isn't 9cscθ, but you are almost there

OpenStudy (anonymous):

use \[1-\sin ^{2}\theta = \cos ^{2}\theta\]

OpenStudy (anonymous):

oh ok...where did the cos come from? I thought sinθ=1/cscθ so in this case it would be 1-sin^2θ=1/cscθ ?

OpenStudy (anonymous):

ok so 9cosθ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Sweet! thanks a bunch!